OpenStudy (anonymous):

How many solutions are there to the following system of equations? -2x + 2y = 8 9x - 9y = 36

OpenStudy (anonymous):

y-x=4........(1) x-y=4.......(2) (1)+(2) 0=4 which is not possible......so no solution

OpenStudy (anonymous):

or you can draw line......we see that they are parallel line.....so .....it has no solution